# Imbalanced Data Clustering using Equilibrium K-Means

### v1

@@ III-C Three Common Smooth Minimum Functions
where $\lambda$ serves as a parameter that controls the degree of approximation, with $\text{LSE}_{\lambda}\to\min$ as $\lambda\to+\infty$. Another common smooth minimum function, called p-Norm, can be defined as $f(x)=x^{p}$, which has the following specific form
@@
and converges to $\min(x_{1},\cdots,x_{K})$ as $p\to+\infty$.
Smooth minimum functions can also be constructed using
@@
A specific example is the Boltzmann operator, which is defined as $g(x)=e^{-\alpha x}$. The Boltzmann operator takes on the form of
@@ III-D1 The relationship between LogSumExp and Maximum-Entropy Fuzzy Clustering
where $\gamma^{(\tau)}_{k}$ is the learning rate at the $\tau$-th iteration. This updating procedure (25) is equivalent to the MEFC algorithm if one set
@@
Such learning rate value is related to the second-order partial derivative of $J_{\text{LSE}}$, which we will discuss later.
@@ III-D2 Towards Lloyd’s algorithm
In the limit $\lambda\to\infty$, $\frac{e^{-\lambda d_{kn}}}{\sum_{i}e^{-\lambda d_{in}}}\to 1$ if $d_{kn}\leq d_{in}\,\forall i\in\{1,\cdots,K\}$, and $\frac{e^{-\lambda d_{kn}}}{\sum_{i}e^{-\lambda d_{in}}}\to 0$ otherwise [^2] We assume $\forall n$, if $k\neq i$, then $d_{kn}\neq d_{in}$.. In this case, the learning rate $\gamma_{k}^{(\tau)}\to 1/N_{k}^{(\tau)}$ where $N_{k}^{(\tau)}$ is the number of data points closest to $\mathbf{c}_{k}^{(\tau)}$, and the updating procedure (25) approaches to Lloyd’s algorithm.
III-D3 Lloyd’s algorithm as a Newton’s method
@@
As $\lambda\to+\infty$, $\partial_{\mathbf{c}_{k}}^{2}{J_{\text{LSE}}(\mathbf{c}_{k}^{(\tau)})}\to N_{k}^{(\tau)}\mathbf{I}$, and the updating procedure (28) asymptotically approaches Lloyd’s algorithm. This indicates that Lloyd’s algorithm is essentially Newton’s method, which aligns with the perspective presented in [29]. There is a close relationship between the learning rate value (26) and the second-order partial derivative of $J_{\text{LSE}}$ (27): the gradient descent with the learning rate of  (26) is equivalent to Newton’s method considering only a portion of the Hessian matrix (ignoring a rank-1 matrix $\mathbf{D}_{kn}$). Consequently, MEFC can be viewed as an approximated Newton’s method. A discussion of the pros and cons of optimizing $J_{\text{LSE}}$ using gradient descent or Newton’s method would be valuable, but it is beyond the scope of this paper. In the following section, we will observe a similar connection between Newton’s method and FKM.
@@ III-D4 The relationship between p-Norm and fuzzy K-means
The objective $J_{\text{PN}}$ has the first partial derivative as
@@
this gradient descent optimization is identical to the FKM algorithm.
The connection between FKM and Newton’s method becomes evident through the second-order partial derivative of $J_{\text{PN}}$, which is
@@
where $\zeta_{kn}=(p+1)\frac{\sum_{i\neq k}d_{in}^{-p}}{d_{kn}\sum_{i}d_{in}^{-p}}$. Hence, FKM can also be viewed as an approximated Newton’s method that ignores the rank-1 matrix $\mathbf{D}_{kn}$.
@@ III-E From Boltzmann Operator to A Novel Clustering Algorithm
Employing the Boltzmann operator to smooth WCSS results in a new objective function, which is
$\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K}}J_{\text{B}}(\mathbf{c}_{1},\cdots,\mathbf{c}_{K})=\sum_{n=1}^{N}\frac{\sum_{i}d_{in}e^{-\alpha d_{in}}}{\sum_{i}e^{-\alpha d_{in}}}.$ (34)
@@
This updating procedure can be reformulated into a two-step iteration procedure, akin to FKM, which entails finding centroids using
@@
where weights $w_{kn}^{(\tau)}$ are given by
$w_{kn}^{(\tau)}=\frac{e^{-\alpha d_{kn}^{(\tau)}}}{\sum_{i}e^{-\alpha d_{in}^{(\tau)}}}[1-\alpha(d_{kn}^{(\tau)}-\frac{\sum_{i}d_{in}^{(\tau)}e^{-\alpha d_{in}^{(\tau)}}}{\sum_{i}e^{-\alpha d_{in}^{(\tau)}}})].$ (39)
The time complexity of one iteration is $O(NK^{2})$. Although the sum of weights for each instance equals one:
$\sum_{k=1}^{K}w_{kn}^{(\tau)}=1,$ (40)
it is worth noting that these weights cannot be interpreted as probabilities nor memberships, since some values are negative. Section III-G will discuss how to define membership in EKM.
@@ III-F Physical Interpretation of Equilibrium K-Means
The second law of thermodynamics asserts that in a closed system with constant external parameters (e.g., volume) and fixed entropy, the internal energy will reach its minimum value at the state of thermal equilibrium. The objective (34) of EKM adheres to this minimum energy principle. This connection can be established by envisioning data points as particles with discrete/quantized energy levels, where the energy value corresponds to the squared Euclidean distance between a data point and a centroid, and the number of energy levels is equivalent to the number of centroids.
According to Boltzmann’s law, at the state of thermodynamic equilibrium, the probability of a particle occupying a specific energy level decreases exponentially with the energy value of that level. Hence, the objective function (34) is equivalent to the expectation of energy for the entire system, and the proposed clustering algorithm seeks centroids that minimize this expectation of energy. Due to this connection, we refer to the proposed algorithm (alternating between (38) and (39)) as equilibrium K-means.
@@ III-G Membership Defined in Equilibrium K-Means
According to the physical interpretation of EKM, the exponential term $e^{-\alpha d_{kn}}$ can be regarded as the unnormalized probability of the $n$-th data point belonging to the $k$-th clusters. Hence, the membership of the $n$-th data point to the $k$-th clusters in EKM can be defined as
@@
Note that, although the membership formula of EKM is the same as that (10) of MEFC, the membership values obtained from EKM and MEFC are typically different because their centroids are calculated using distinct formulas. This paper does not intend to compare the memberships in different fuzzy clustering algorithms; therefore, we only define membership in EKM without further discussion.
@@ III-H Convergence of Smooth K-Means
HKM, FKM, MEFC, and EKM can be viewed as special cases of SKM, which can be generalized as the following gradient descent algorithm:
@@
The convergence condition of this gradient descent algorithm is given as follows:
@@
It can be easily verified that when the smooth minimum function $h$ is LogSumExp (19) or p-Norm (20), the above convergence condition can be satisfied with any initial centroids. Although the Boltzmann operator (22) is not concave, it is a smooth approximation of a concave function (the minimum function). Hence, EKM also demonstrates good convergence behavior, which has been empirically verified in our experiments.
@@ IV-A Case Study
This section explores and compares the behavior of different KM algorithms by examining their new objective functions in some examples. We select these examples because their data structures are well-designed, making it easier to illustrate the clustering characteristics.
We generate one-dimensional datasets comprising two classes of data points by sampling from two normal distributions, drawing $N_{1}$ samples from a distribution with mean $\mu_{1}$ and unit variance, and $N_{2}$ samples from another with mean $\mu_{2}$ and unit variance. Using different parameter combinations, we create four datasets: 1. A balanced, non-overlapping dataset ($N_{1}=N_{2}=50$, $\mu_{1}=-5$, and $\mu_{2}=+5$; Fig. 2a); 2. A balanced, overlapping dataset ($N_{1}=N_{2}=50$, $\mu_{1}=-0.5$, and $\mu_{2}=+0.5$; Fig. 2b); 3. An imbalanced, non-overlapping dataset ($N_{1}=5000$, $N_{2}=50$, $\mu_{1}=-5$, and $\mu_{2}=+5$; Fig. 2c); 4. An imbalanced, compact dataset ($N_{1}=2000$, $N_{2}=50$, $\mu_{1}=-2$, and $\mu_{2}=+2$; Fig. 2d).
We fix the first centroid at $\mu_{1}$, and plot the objectives of HKM (WCSS (12)), FKM ($J_{\text{PN}}$ (29)), MEFC ($J_{\text{LSE}}$ (23)), and EKM ($J_{\text{B}}$ (34)) as a function of the position of the second centroid in Fig. 2. Fig. 2a demonstrates the similarity in objective functions for the four clustering algorithms, indicating comparable performance on a balanced dataset. Fig. 2b reveals that FKM and MEFC are better suited for overlapped data, as they are less prone to a local minimum. In Fig. 2c and Fig. 2d, data imbalance cases show that the local and global minimum points of the objective functions of HKM, FKM, and MEFC are biased towards the center of the large class, i.e., $\mu_{1}$. In contrast, EKM does not have an obvious local minimum and its global minimum point aligns with the correct cluster center, i.e., $\mu_{2}$, highlighting its superiority in handling imbalanced data.
Figure 2: Objectives of four KM algorithms as a function of the first centroid fixed at $\mu_{1}$ and the second centroid moving along the $x$-axis. The yellow, purple, green, and blue curves are the objective functions of HKM, FKM, MEFC, and EKM, respectively. (a), (b), (c), and (d) present objective functions for different data distributions.
@@ IV-B The Analysis of EKM’s Effectiveness on Imbalanced Data
We demonstrate the effectiveness of EKM by analyzing the gradient of its objective. The gradient of the smoothed WCSS with respect to the distance between data points and centroids can be interpreted as the force exerted by a spring. Positive gradient values represent attractive forces, while negative values represent repulsive forces. The gradient is also associated with membership, where the membership (7) of FKM is equivalent to a power exponent ($1/m$) of $\partial J_{\text{PN}}/\partial d_{kn}$, and the membership (10) of MEFC is equivalent to $\partial J_{\text{LSE}}/\partial d_{kn}$.
A simple example is given to plot the gradient values of different smoothed objectives. We fix two centroids at $-1$ and $+1$, respectively, and move a data point along the $x$-axis. In Fig. 3a we display the gradients of WCSS, $J_{\text{LSE}}$, $J_{\text{PN}}$, and $J_{\text{B}}$ with respect to the distance between the data point and the second centroid. As evident from the figure, data points on the side of the first centroid do not impact the second centroid of HKM, but they do attract the second centroids of FKM and MEFC. This finding supports the previous result in [31] that FKM has a stronger uniform effect than HKM. However, for EKM, data points near the first centroid have repulsive forces on the second centroid, which compensates for the attraction of points on the side of the first centroid to the second centroid, effectively reducing the uniform effect. Furthermore, it is worth noting that for EKM, data points near the second centroid have the strongest attraction forces. Thus, the second centroid is anchored by these data points, making its position less susceptible to noise or outliers. This is the advantage of EKM over imbalanced data clustering algorithms that assign weights to data points based on cluster size.
Figure 3: Gradients of four KM algorithms as a function of a data point moving along the $x$-axis. (a) Gradients of HKM, FKM, MEFC, and EKM with fixed smoothness parameters. (b) Gradients of FKM. (c) Gradients of MEFC. (d) Gradients of EKM.
@@ IV-C The Choice of $\alpha$
The smoothing parameter $\alpha$ has an impact on the clustering result in EKM, but the optimal choice remains unknown, which is not a difficulty unique to EKM. The FKM algorithm also struggles with selecting the optimal fuzzifier value, $m$. Despite numerous studies discussing the selection of $m$, a widely accepted solution has yet to be found [32].
Determining optimal parameters for a clustering algorithm often involves trial and error, as the real data structure is unknown. As a rule of thumb, when the data space dimension is less than or equal to three, setting $\alpha=1$ appears effective after normalizing the data to have a zero mean and unit variance for each dimension. As the data space dimension increases, the data-centroid distance $d_{kn}$ tends to grow, necessitating a decrease in $\alpha$ to ensure that the exponential term $e^{-\alpha d_{kn}}$ remains within a normal range. In this case, we suggest setting $\alpha$ to ten times the data variance and gradually reducing it. A proper $\alpha$ value indicates the presence of repulsive forces, resulting in a sudden increase in centroid distances. Ultimately, techniques for visualizing clustering results can be employed to fine-tune the $\alpha$ value. This strategy is proven effective by experimental examples of deep clustering on the MNIST dataset in Section VI. We further discuss and validate the choice of $\alpha$ in the supplemental material.

### v2

@@ III-C Three Common Smooth Minimum Functions
where $\lambda$ is a parameter controlling the degree of approximation, with $\text{LSE}_{\lambda}\to\min$ as $\lambda\to+\infty$.
Define $f(x)=x^{p}$, we can have another common smooth minimum function, called p-Norm, which has the following specific form
@@
and converges to $\min(x_{1},\cdots,x_{K})$ as $p\to+\infty$. An additional definition is necessary for the domain of the p-Norm function to legally include zeros: $\text{PN}_{p}(x_{1},\cdots,x_{K})=0$ if some $x_{i}=0$ where $1\leq i\leq K$.
Smooth minimum functions can also be constructed by
@@
A specific example is the Boltzmann operator, where $g(x)=e^{-\alpha x}$. The Boltzmann operator takes on the form of
@@ III-D1 The relationship between LogSumExp and Maximum-Entropy Fuzzy Clustering
where $\gamma^{(\tau)}_{k}$ is the learning rate at the $\tau$-th iteration. This updating procedure (25) is equivalent [^2] It should be noted that the equivalence mentioned in this paper is in the sense of the algorithm level, not in the criterion level. to the MEFC algorithm if one set
@@
Such learning rate value is related to the second-order partial derivative of $J_{\text{LSE}}$, which we will discuss later. The membership (10) of MEFC is identical to $\partial J_{\text{LSE}}/\partial d_{kn}$.
@@ III-D2 Towards Lloyd’s algorithm
In the limit of $\lambda\to\infty$, $\frac{e^{-\lambda d_{kn}}}{\sum_{i}e^{-\lambda d_{in}}}\to 1$ if $d_{kn}\leq d_{in}\,\forall i\in\{1,\cdots,K\}$, and $\frac{e^{-\lambda d_{kn}}}{\sum_{i}e^{-\lambda d_{in}}}\to 0$ otherwise [^3] We assume $\forall n$, if $k\neq i$, then $d_{kn}\neq d_{in}$.. In this case, the learning rate $\gamma_{k}^{(\tau)}\to 1/N_{k}^{(\tau)}$ where $N_{k}^{(\tau)}$ is the number of data points closest to $\mathbf{c}_{k}^{(\tau)}$, and the updating procedure (25) approaches to Lloyd’s algorithm.
III-D3 Lloyd’s algorithm and Newton’s method
@@
As $\lambda\to+\infty$, $\partial_{\mathbf{c}_{k}}^{2}{J_{\text{LSE}}(\mathbf{c}_{k}^{(\tau)})}\to N_{k}^{(\tau)}\mathbf{I}$, and the updating procedure (28) approaches Lloyd’s algorithm. This indicates that Lloyd’s algorithm is essentially Newton’s method, which aligns with the perspective presented in [28]. There is a close relationship between the learning rate value (26) and the second-order partial derivative of $J_{\text{LSE}}$ (27): the gradient descent with the learning rate of  (26) is equivalent to an approximated Newton’s method in the sense that one term (the rank-1 matrix $\mathbf{D}_{kn}$) in (27) is ignored. A discussion of the pros and cons of optimizing $J_{\text{LSE}}$ using gradient descent or Newton’s method would be valuable, but it is beyond the scope of this paper. In the following section, we will observe the same connection between Newton’s method and FKM.
@@ III-D4 The relationship between p-Norm and fuzzy K-means
The objective $J_{\text{PN}}$ has the first partial derivative of
@@
this gradient descent iteration is equivalent to the FKM algorithm. The membership (7) of FKM is equivalent to a power exponent ($1/m$) of $\partial J_{\text{PN}}/\partial d_{kn}$.
The connection between FKM and Newton’s method becomes evident by taking the second-order partial derivative of $J_{\text{PN}}$, which is
@@
where $\zeta_{kn}=(p+1)\frac{\sum_{i\neq k}d_{in}^{-p}}{d_{kn}\sum_{i}d_{in}^{-p}}$. Hence, FKM can also be viewed as an approximated Newton’s method that ignores the term of the rank-1 matrix $\mathbf{D}_{kn}$.
@@ III-E From Boltzmann Operator to A Novel Clustering Algorithm
Employing the Boltzmann operator to smooth WCSS results in
$\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K}}J_{\text{B}}(\mathbf{c}_{1},\cdots,\mathbf{c}_{K})=\sum_{n=1}^{N}\frac{\sum_{i=1}^{K}d_{in}e^{-\alpha d_{in}}}{\sum_{i=1}^{K}e^{-\alpha d_{in}}}.$ (34)
@@
This updating procedure can be reformulated into a two-step iteration procedure akin to FKM, that is
@@
where weights $w_{kn}^{(\tau)}$ are calculated by
$w_{kn}^{(\tau)}=\frac{e^{-\alpha d_{kn}^{(\tau)}}}{\sum_{i=1}^{K}e^{-\alpha d_{in}^{(\tau)}}}[1-\alpha(d_{kn}^{(\tau)}-\frac{\sum_{i=1}^{K}d_{in}^{(\tau)}e^{-\alpha d_{in}^{(\tau)}}}{\sum_{i=1}^{K}e^{-\alpha d_{in}^{(\tau)}}})].$ (39)
The time complexity of one iteration is $O(NK^{2})$. Although the sum of weights equals one, i.e.,
$\sum_{k=1}^{K}w_{kn}^{(\tau)}=1,\,\forall n,$ (40)
it is worth noting that these weights cannot be interpreted as probabilities or memberships since some weight values are negative. The membership of EKM is defined in Section III-G.
@@ III-F Physical Interpretation of Equilibrium K-Means
The second law of thermodynamics asserts that in a closed system with constant external parameters (e.g., volume) and fixed entropy, the internal energy will reach its minimum value at the state of thermal equilibrium. The objective (34) of EKM follows this minimum energy principle. This connection can be established by envisioning data points as particles with discrete/quantized energy levels, where the number of energy levels is equivalent to the number of centroids, and the energy value corresponds to the squared Euclidean distance between a data point and a centroid.
Boltzmann’s law tells that at the state of thermodynamic equilibrium, the probability of a particle occupying a specific energy level decreases exponentially with the increase of the energy value of that level. Hence, the objective function (34) equals the expectation of the entire system’s energy, and EKM seeks centroids to minimize this energy expectation. Due to this connection, we refer to the proposed algorithm (alternating between (38) and (39)) as equilibrium K-means.
@@ III-G Membership Defined in Equilibrium K-Means
According to the physical interpretation of EKM, the exponential term $e^{-\alpha d_{kn}}$ can be interpreted as the unnormalized probability of the $n$-th data point belonging to the $k$-th clusters. Hence, the membership of the $n$-th data point to the $k$-th cluster can be defined as
@@
Note that, although the membership formula of EKM is the same as that (10) of MEFC, the values are different because centroids are calculated using distinct formulas. We have no intention in this paper to compare membership defined in different fuzzy clustering algorithms; therefore, we only define membership in EKM and leave the further discussion to subsequent research.
@@ III-H Convergence of Smooth K-Means
HKM, FKM, MEFC, and EKM are special cases of SKM, which can be generalized as the following gradient descent algorithm:
@@
Different KM algorithms can be obtained by taking $h$ the corresponding explicit form (refer to Section III-D to III-E). This general form facilitates the uniform study of KM algorithms. Below we give a convergence guarantee conditioning on the properties of $h$:
@@
It can be easily verified that when the smooth minimum function $h$ is LogSumExp (19) and p-Norm (20) in which case SKM (42) is equivalent to MEFC and FKM, respectively, the above convergence condition can be satisfied with any initial centroids. Although the Boltzmann operator (22) is not concave, it is a smooth approximation of a concave function (the minimum function). Hence, EKM also exhibits good convergence behavior in numerical experiments.
@@ IV-A Case Study
This section presents an empirical analysis of the behavior of different KM algorithms by examining their reformulated objective functions in some examples with well-designed data structures. Datasets comprising two classes of one-dimensional data points are generated by sampling from two normal distributions, drawing $N_{1}$ samples from a distribution with mean $\mu_{1}$ and unit variance, and $N_{2}$ samples from another with mean $\mu_{2}$ and unit variance. Using different parameter combinations, we generate four datasets: 1. A balanced, non-overlapping dataset ($N_{1}=N_{2}=50$, $\mu_{1}=-5$, and $\mu_{2}=+5$; Fig. 2a); 2. A balanced, overlapping dataset ($N_{1}=N_{2}=50$, $\mu_{1}=-0.5$, and $\mu_{2}=+0.5$; Fig. 2b); 3. An imbalanced, non-overlapping dataset ($N_{1}=5000$, $N_{2}=50$, $\mu_{1}=-5$, and $\mu_{2}=+5$; Fig. 2c); 4. An imbalanced, overlapping dataset ($N_{1}=2000$, $N_{2}=50$, $\mu_{1}=-2$, and $\mu_{2}=+2$; Fig. 2d).
We fix the first centroid at $\mu_{1}$, and plot the reformulated objectives of HKM (WCSS (12)), FKM ($J_{\text{PN}}$ (29)), MEFC ($J_{\text{LSE}}$ (23)), and EKM ($J_{\text{B}}$ (34)) as a function of the position of the second centroid in Fig. 2. The four objective functions behave similarly on the first two balanced datasets, but it is worth noting that the last two imbalanced datasets. Fig. 2c and Fig. 2d show that the local and global minimum points of the objective functions of HKM, FKM, and MEFC are biased towards the center of the large cluster, i.e., $\mu_{1}$. In contrast, EKM does not have an obvious local minimum and its global minimum point aligns with the true cluster center, i.e., $\mu_{2}$, highlighting EKM’s superiority in handling imbalanced data.
Figure 2: Objectives as a function of the second centroid moving along the $x$-axis (the first centroid is fixed at $\mu_{1}$). The yellow, purple, green, and blue curves are reformulated objective functions of HKM, FKM, MEFC, and EKM, respectively. (a), (b), (c), and (d) present objective functions under different data distributions.
@@ IV-B The Analysis of EKM’s Effectiveness on Imbalanced Data
We analyze the effectiveness of EKM on imbalanced data based on the gradient of its objective. The gradient of the smoothed WCSS with respect to the data-centroid distance (i.e., $\partial J/\partial d_{kn}$) can be interpreted as the force exerted by a spring. A positive gradient value represents an attractive force, while a negative value represents a repulsive force.
A simple example is given to plot the gradient values of different KM’s objectives. We fix two centroids at $-1$ and $+1$, respectively, and move a data point along the $x$-axis. In Fig. 3a we display the gradients of WCSS, $J_{\text{LSE}}$, $J_{\text{PN}}$, and $J_{\text{B}}$ with respect to the distance between the data point and the second centroid. As evident from the figure, data points on the side of the first centroid do not impact the second centroid of HKM, but they do attract the second centroids of FKM and MEFC. This finding supports the statement in [30] that FKM has a stronger uniform effect than HKM. On the other hand, data points near the first centroid have repulsive forces on the second centroid of EKM, which compensate for the attraction from other data points, effectively reducing the uniform effect. It is worth noting that data points near the second centroid of EKM have the strongest attraction forces to this centroid. Consequently, EKM’s centroids are anchored by their surrounding data points, making their position less influenced by noise and outliers.
Figure 3: Gradients as a function of a data point moving along the $x$-axis. (a) Gradients of WCSS, $J_{\text{PN}}$, $J_{\text{LSE}}$, and $J_{\text{B}}$ with fixed smoothing parameters. (b), (c), and (d) are gradients of $J_{\text{PN}}$, $J_{\text{LSE}}$, and $J_{\text{B}}$ with varied smoothing parameters, respectively.
@@ IV-C The Choice of $\alpha$
The smoothing parameter $\alpha$ impacts the performance of EKM, but the optimal choice remains unknown. This is not a difficulty unique to EKM. The FKM algorithm also struggles with selecting the optimal fuzzifier value, $m$. Despite numerous studies discussing the selection of $m$, a widely accepted solution has yet to be found [31].
As a rule of thumb, when the dimension of the data space is less than or equal to three, setting $\alpha=1$ appears effective after normalizing the data to have a zero mean and unit variance for each dimension. As the data space dimension increases, the data-centroid distance $d_{kn}$ increases, necessitating a decrease in $\alpha$ to ensure that the exponential term $e^{-\alpha d_{kn}}$ falls within a normal range. Hence, in the case of a higher dimensional data clustering, we suggest initially setting $\alpha$ to ten times the data variance and gradually reducing it until a sudden increase in centroid-centroid distance is observed. Because the increase in centroid distance implies the emergence of repulsive forces that reduce the uniform effect. This strategy is applied in deep clustering on the MNIST dataset in Section VI. An adaptive selection formula of $\alpha$ and its validation are given in the supplemental material.